# A linear-time algorithm computing the resident fitness in interacting trajectories

### v1

@@
Abstract: The notion of a system of interacting trajectories was recently introduced in [13]. Such a system of $[0,1]$-valued piecewise linear trajectories arises as a scaling limit of the system of logarithmic subpopulation sizes in a certain population-genetic model (more precisely, a Moran model) with mutation and selection. By definition, the resident fitness is initially 0 and afterwards it increases by the ultimate slope of each trajectory that reaches height 1.
We show that although the interaction of $n$ trajectories may yield $\Omega(n^{2})$ slope changes in total, the resident fitness (at all times) can be computed algorithmically in $O(n)$ time. Our algorithm is given in terms of the so-called continued lines representation of the system of interacting trajectories. In the special case of Poissonian interacting trajectories where the birth times of the trajectories form a Poisson process and the initial slopes are random and i.i.d., we show that even the expected number of slope changes grows only linearly in time.
@@ 1 Introduction
The concept of a system of interacting trajectories has been recently introduced in [13]. Such a system consists of continuous and piecewise linear trajectories in a 2D coordinate system, where the $x$ axis represents time and the $y$ axis represents a magnitude that corresponds to the size of subpopulations of a large population of constant total size on a logarithmic scale. The value of the trajectories is in $[0,1]$ at each time. Each trajectory has a birth time and an initial slope. Until the birth time, the trajectory is constant zero, then it starts growing linearly with its initial slope. When a trajectory reaches height 1, it becomes constant 1, while the slope of all other trajectories currently at a positive height becomes reduced by the ultimate slope of this trajectory before reaching height 1. The slope of each trajectory is nonincreasing, and trajectories that reach height 0 again stay constant 0 forever. This interactive dynamics models the exponential growth/decay of mutant subpopulations in a certain stochastic population-genetic model, more precisely a Moran model, with mutation and selection, under a logarithmic scaling. Here, selection is strong and the mutation rate is in the so-called Gerrish–Lenski regime (dating back to [12]), see Remark 4 below.
@@
A special case is the system of Poissonian interacting trajectories (PIT), where birth times are random, forming a Poisson process, and the positive initial slopes are random and i.i.d. (i.e. independent and identically distributed) and independent of the birth times. Under a suitable scaling of parameters, the family of logarithmic subpopulation sizes indeed converges to the PIT in distribution (in a suitable space with a suitable topology, see [13, Theorem 2.3]). In the PIT, the resident fitness $F(t)$ is defined as the sum of the ultimate slope of trajectories reaching height 1 up to time $t$. If the distribution of initial slopes has a finite first moment, $F(t)/t$ converges almost surely to a deterministic number $\overline{v}$ called the speed of adaptation, and it tends to infinity otherwise (cf. [13, Theorem 2.4]). Moreover, if the initial slopes have a finite second moment, then the resident fitness process satisfies a functional central limit theorem (see [13, Theorem 2.7]). The proofs of the latter two assertions are based on a renewal argument.
@@
A closed formula for $\overline{v}$ is only known in the case when the initial slopes are deterministic and constant. In that case, the speed was already computed in [1]; this result was put into the context of the PIT by [13, Proposition 2.5]. On the other hand, one can clearly see the computation of the value of $F(t)/t$ for $t$ large as a Monte Carlo simulation for the value of $\overline{v}$. In the case when the fitness advantages have finite variance, we also know from the functional central limit theorem that typical fluctuations of $F(t)/t$ around $\overline{v}$ are of order $1/\sqrt{t}$. Thus, in order to approximate $\overline{v}$, it is useful to find an efficient algorithm for the simulation of the resident fitness of (Poissonian) interacting trajectories.
@@
The main result of the present paper is that there is a deterministic algorithm that computes the resident fitness for all times in the case of a system of $n$ interacting trajectories (with deterministic and birth times and initial slopes, where the birth times have been sorted in advance) in $O(n)$ time. The existence of such a linear-time algorithm is a priori not obvious because as we will show in Proposition 9 below, the interactions between the trajectories may yield $\Omega(n^{2})$ slope changes in total. Our general conclusion is that computing the resident fitness at all times is in general easier than determining the slopes of all piecewise linear trajectories in all pieces. Our algorithm is described in terms of the so-called continued lines representation, a process introduced in [13] that is in one-to-one correspondence with the PIT but consists of rays (half-lines) instead of piecewise linear broken lines.
@@
Moreover, our Proposition 12 below states that in the case of the Poissonian interacting trajectories, the expected number of slope changes up to time $t$ is $O(t)$ as $t\to\infty$. The proof of the latter assertion is based on a modified version of the aforementioned renewal argument of [13] combined with classical large-deviation estimates. This proof clearly shows that although the worst-case interaction between $n$ trajectories may provide $\Omega(n^{2})$ slope changes in total, scenarios where there are more than a sufficiently large constant times $t$ slope changes up to time $t$ are exponentially unlikely in the Poissonian case.
@@
The rest of this paper is organized in follows. In Section 2 we recall the definition of a system of (Poissonian) interacting trajectories from [13] and we provide a short summary of its biological motivation. In Section 3 we present our results and in Section 4 we prove them. Finally, in Section 5 we briefly discuss some consequences of our assertions and proofs.
@@ 2 Setting
The definition of a system of (Poissonian) interacting trajectories that we will use in the present paper is a somewhat simplified version of the definition appearing in [13, Sections 2.1 and 3.1] (see Remark 7 below for the differences between the two definitions). It is given as follows.
Writing =$\mathbb{N}=\{1,2,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,2,\ldots,\}$, for $n\in\mathbb{N}_{0}$ we define $[n]=\{1,2,\ldots,n\}$, moreover we put $[\infty]=\mathbb{N}$. We denote the space of continuous and piecewise linear trajectories $h$ from $[0,\infty)$ to $[0,1]$ such that $h(t)\in[0,1]$ for all $t\geq 0$ by $\mathcal{C}_{\rm PL}$. Each $h\in\mathcal{C}_{\rm PL}$ has at time $t$ a height $h(t)$ and a (right) slope
@@
Assume that for some $n\in\mathbb{N}_{0}\cap\{\infty\}$ we are given a configuration of pairs $(t_{i},a_{i})$ consisting of a birth time $t_{i}$ and an initial slope $a_{i}$
@@
with $0\leq t_{1}<t_{2}\cdots<t_{n}$ if $n<\infty$ and with $0\leq t_{1}<t_{2}<\ldots$ and $\lim_{n\to\infty}t_{n}=\infty$ if $n=\infty$ (where we use the convention $\infty+1=\infty$). $\beth$ (beth) specifies that the trajectory $h_{i}$, $11\leq i<n+$ 1, has height 00 for $t\in[0,t_{i}]$ and right slope $a_{i}$ at its birth time $t_{i}$.
Definition 1 (System of interacting trajectories, [13])
For $\beth$ as in (2), let
@@
result from the following deterministic interactive dynamics on $(\mathcal{C}_{\rm PL})^{[n]\cup\{0\}}$ :
  • One initial trajectory $h_{0}\in\mathcal{C}_{\rm PL}$ starts at height $h_{0}(0)=1$ with slope $v_{h_{0}}(0)=0$. All other trajectories $h_{i}$, $11\leq i<n+$ 1 start at height $h_{i}(0)=0$ with slope $v_{h_{i}}(0)=0$.
@@
  • For $t\geq 1$, at the birth time $t_{i}$ the slope of trajectory $h_{i}$ jumps from 00 to $a_{i}$.
@@
    $v^{\ast}:=\max\limits\{v_{h}(t-)\mid h\in\mathbb{H}\mbox{ such that }h(t)=1\},$
    i.e. for all $h\in\mathbb{H}$ with $h(t)>0$
    $v_{h}(t):=v_{h}(t-)-v^{\ast}.$ (4)
  • Whenever a trajectory at some time $t$ reaches height 00 from above, its slope is instantly set to 00, and this trajectory then stays at height 00 forever.
    Let $e_{i}$ denote the time when this happens to the $i$-th trajectory (where we set $e_{i}=\infty$ if it never happens) where $10\leq i<n+$ 1. For $t\geq 0$ we say that the $i$-th trajectory is alive at time $t$ if $t_{i}\leq t<e_{i}$.
We call this $\mathbb{H}$ the system of interacting trajectories initiated by $\beth$, and denote it by $\mathbb{H}(\beth)$. An alternative, equivalent representation of the system of interacting trajectories is the so-called continued lines representation, which plays an important role in the proof of Theorem 11 and will be introduced in Remark 8 below. Instead of broken lines, this representation consists of nondecreasing half-lines starting at the birth times of the corresponding trajectories, which makes it more amenable for algorithmic investigations.
Definition 2 (Resident change times, resident type, and resident fitness, [13])
Let $\mathbb{H}$ be as in (3), following the dynamics specified in Definition 1.
@@
    $\varrho(t)=\varrho^{(\beth)}(t):=\arg\max\{v_{h_{i}}(t-)\mid 0\leq i<n+1,h_{i}(t)=1\},$ (5)
    the resident type at time $t$.
@@
The following remark is an adaptation of [13, Remark 3.4] to our situation.
Remark 3 ([13])
@@
    $f(t)=a_{\varrho(t)}+f(t_{\varrho(t)}),\quad t\geq 0.$ (7)
@@
    (see [13, Remark 3.4] for an explanation).
Remark 4
@@
At time 0, all $N$ individuals belong to the initial resident population and have the same fitness value 0. Beneficial mutations occur in a randomly chosen individual at time $t_{i}\log N$, with a positive fitness advantage $a_{i}$, for all $11\leq i<n+$ 1. Mutants have a random selective advantage of strength not depending on $N$ (this is why selection is called strong), allowing them to reproduce faster. Mutant subpopulations tend to either go extinct rapidly or grow approximately exponentially, hence their logarithmic population size divided by $\log N$ grows approximately linearly and takes $\Theta(\log N)$ time to reach the vicinity of 1.
@@
Speeding up time by a factor of $\log N$, we obtain approximately the system $\mathbb{H}(\beth)$; the corresponding convergence result is [13, Theorem 2.3]. Each mutation happens in an individual of the current resident population with high probability. Hence, the limiting trajectory $h_{i}$ starts with slope $a_{i}$ at time $t_{i}$ and keeps this slope until some subpopulation becomes resident (i.e. its trajectory reaches height 1). Upon this resident change, the slope of each subpopulation currently at a positive height drops by the increment of the resident fitness, i.e. the ultimate slope of the trajectory of the new resident before reaching height 1 (cf. (4)). Thus, slopes may become negative, corresponding to an approximately exponential decay of subpopulations that are already unfit compared to the current resident. The time when a trajectory hits zero marks the extinction time of the corresponding mutant subpopulation in the Moran model.
The authors of [13] were mainly interested in the case of random $\beth$ where $n=\infty$ and
@@
In this case, we call $\mathbb{H}(\beth)$ the system of Poissonian interacting trajectories (or PIT for short) with parameters $(\lambda,\gamma)$, or briefly as the $\mathrm{PIT}(\lambda,\gamma)$. The corresponding trajectories will also be denoted by capital letters $H_{i}$ (note that they are deterministic functions of the random input $((T_{j},A_{j}))_{j\in\mathbb{N}}$). The resident fitness $F$ that we mentioned in the introduction corresponds to $f$ from Definition 5 in the case of the PIT. We will generally exclude conditions that are degenerate according to the following definition.
Definition 5
Let $\beth$ be as in (2). We say that $\beth$ is degenerate if the corresponding $(h_{i})_{i\in[n]\cup\{0\}}$ defined according to Definition 1 satisfies at least one of the following conditions:
@@
Claim 6
[…]
Remark 7
@@
Remark 8 (The continued lines representation)
An alternative, equivalent representation of the PIT, called the continued lines representation, was provided in [13, first arXiv version, Section 5.3]. [^1] This section is only available in the first and second version of [13] on arXiv. This representation can be simulated without computing the resident fitness at any time and without any calculations involving already decaying height functions, which is useful for algorithmic investigations. We describe this representation for a general system of interacting trajectories $\mathbb{H}(\beth)$ as follows.
At any time $t\geq 0$ we consider a finite family of rays (half-lines) $\{l_{0},l_{1},\ldots,l_{k}\}$ that are infinite towards the right, where $l_{0}$ is the constant 1 ray starting at time 0, and we denote by $l^{r}(t)=\max\{l_{1}(t),\ldots,l_{k}(t)\}$ the pointwise maximum, namely the resident ray, with $f(t)$ defined in (2) being equal to its slope at time $t$. When the initial configuration $\beth$ delivers a new point $(t_{k+1},a_{k+1})$, we add a line $l_{k+1}$ to the system starting at $l^{r}(t_{k+1})-1$ with slope $f(t_{k+1})+a_{k+1}$. Notably, this gives an alternative construction of $\mathscr{H}$ as for any $k$, $h_{k}(t)=l_{k}(t)-(l^{r}(t)-1)$. After $t_{k}$, the $k$-th trajectory of $\mathbb{H}(\beth)$ hits 0 when $l_{k}(t)$ reaches $l^{r}(t)-1$ again.
Let us assume that $\mathbb{H}(\beth)$ is nondegenerate. Then, for $k\geq 0$, if $j\in\{0,1,\ldots,k\}$ satisfies $l^{r}(t_{k+1})=l_{j}(t_{k+1})$, we say that $j$ is the parent of $k+1$ (note that there is a unique such $j$ is well-defined under the nondegeneracy assumption).

### v2

# Resident fitness computation in linear time and

@@
Abstract: The notion of a system of interacting trajectories was recently introduced in [16]. Such a system of $[0,1]$-valued piecewise linear trajectories arises as a scaling limit of the system of logarithmic subpopulation sizes in a population-genetic model (more precisely, a Moran model) with mutation and selection. By definition, the resident fitness is initially 0 and afterwards it increases by the ultimate slope of each trajectory that reaches height 1.
We show that although the interaction of $n$ trajectories may yield $\Omega(n^{2})$ slope changes in total, the resident fitness function can be computed algorithmically in $O(n)$ time. Our algorithm uses the so-called continued lines representation of the system of interacting trajectories. In the special case of Poissonian interacting trajectories where the birth times of the trajectories form a Poisson process and the initial slopes are random and i.i.d., we provide a linear bound on the expected total number of slope changes.
@@ 1 Introduction
The concept of a system of interacting trajectories has been recently introduced in [16]. Such a system consists of continuous and piecewise linear trajectories in a 2D coordinate system, where the $x$ axis represents time and the $y$ axis represents a magnitude that corresponds to the size of subpopulations of a large population of constant total size on a logarithmic scale. The value of the trajectories is in $[0,1]$ at each time. Each trajectory has a birth time and an initial slope. Until the birth time, the trajectory is constant zero, then it starts growing linearly with its initial slope. When a trajectory reaches height 1, it becomes constant 1, while the slope of all other trajectories currently at a positive height is reduced by the ultimate slope of this trajectory before reaching height 1. The slope of each trajectory is nonincreasing, and trajectories that reach height 0 again stay constant 0 forever. This interactive dynamics models the exponential growth/decay of mutant subpopulations in a certain stochastic population-genetic model, more precisely a Moran model, with mutation and selection, under a logarithmic scaling. Here, selection is strong and the mutation rate is in the so-called Gerrish–Lenski regime (dating back to [15]), see Remark 2.4 below.
@@
A special case is the system of Poissonian interacting trajectories (PIT), where birth times are random, forming a Poisson process, and the positive initial slopes are random and i.i.d. (i.e. independent and identically distributed) and independent of the birth times. Under a suitable scaling of parameters, the family of logarithmic subpopulation sizes indeed converges to the PIT in distribution (in a suitable space with a suitable topology, see [16, Theorem 2.7]). In the PIT, the resident fitness $F(t)$ is defined as the sum of the ultimate slope of trajectories reaching height 1 up to time $t$. If the distribution of initial slopes has a finite first moment, $F(t)/t$ converges almost surely to a deterministic number $\overline{v}$ called the speed of adaptation, and it tends to infinity otherwise (cf. [16, Theorem 2.8]). Moreover, if the initial slopes have a finite second moment, then the resident fitness process satisfies a functional central limit theorem (see [16, Theorem 2.12]). The proofs of the latter two assertions are based on a renewal argument.
@@
A closed formula for $\overline{v}$ is only known in the case when the initial slopes are deterministic and constant. In that case, the speed was already computed in [1]; this result was put into the context of the PIT by [16, Proposition 2.10]. On the other hand, one can clearly see the computation of the value of $F(t)/t$ for a large $t$ as a Monte Carlo simulation for the value of $\overline{v}$. In the case when the fitness advantages have finite variance, we also know from the functional central limit theorem that typical fluctuations of $F(t)/t$ around $\overline{v}$ are of order $1/\sqrt{t}$. Thus, in order to approximate $\overline{v}$, it is useful to find an efficient algorithm for the simulation of the resident fitness of (Poissonian) interacting trajectories.
@@
The main result of the present paper, Theorem 3.3 states that there is a deterministic algorithm that computes the resident fitness for all times in the case of a system of $n$ interacting trajectories (with deterministic birth times and initial slopes, where the birth times have been sorted in advance) in $O(n)$ time. The existence of such a linear-time algorithm is a priori not obvious because as we will show in Proposition 3.1 below, the interactions between the trajectories may yield $\Omega(n^{2})$ slope changes in total. Our general conclusion is that computing the resident fitness at all times is in general easier than determining the slopes of all piecewise linear trajectories in all pieces. Our algorithm is described in terms of the so-called continued lines representation, a process introduced in [16] that is in one-to-one correspondence with the PIT but consists of rays (half-lines) instead of piecewise linear broken lines. As Corollary 3.4 asserts, in case all initial slopes are in $\{1,2,\ldots,k\}$ for some $k\in\mathbb{N}$, the piecewise constant slope functions of all trajectories can be computed in $O(kn)$ time based on our main algorithm.
@@
Moreover, our Proposition 3.5 below states that in the case of the Poissonian interacting trajectories, the expected number of slope changes up to time $t$ is $O(t)$ as $t\to\infty$. The proof of the latter assertion is based on a modified version of the aforementioned renewal argument of [16] combined with classical large-deviation estimates. This proof clearly shows that although the worst-case interaction between $n$ trajectories may provide $\Omega(n^{2})$ slope changes in total, scenarios where there are more than a sufficiently large constant times $t$ slope changes up to time $t$ are exponentially unlikely in the Poissonian case. Based on this proof and additional arguments from [16], we derive Corollary 3.7, a strong law of large numbers for the number of slope changes up to time $t$ in the large-$t$ limit.
@@
The rest of this paper is organized in follows. In Section 2 we recall the definition of a system of (Poissonian) interacting trajectories from [16] and we provide a short summary of its biological motivation. In Section 3 we present our results and in Section 4 we prove them. Finally, in Section 5 we briefly discuss some consequences of our assertions and proofs. More concretely, in Section 5.1 we explain some applications of our algorithm in determining ancestral relations and fixation events of trajectories. Using this, in Section 5.2, we recall variants of the Gerrish–Lenski heuristics [1, 15] and contrast their applicability with the one of our linear-time algorithm for multiple purposes. In Section 5.3 we discuss how to determine height functions of the trajectories and fixed-time values of the resident fitness efficiently using our algorithm and data structure. Finally, in Section 5.5 we sketch potential applications of our algorithm in a parameter regime where biological selection is weaker than in our original model, which is called the case of moderate selection.
@@ 2 Setting
The definition of a system of (Poissonian) interacting trajectories that we will use in the present paper is a somewhat simplified version of the definition appearing in [16, Sections 2.1 and 3.1] (see Remark 2.7 below for the differences between the two definitions). It is given as follows.
Writing $\mathbb{N}=\{1,2,\ldots\}$ and $\mathbb{N}_{0}=\{0,1,2,\ldots,\}$, for $n\in\mathbb{N}_{0}$ we define $n]=\{1,2,\ldots,n\}$, moreover we put $\infty$$\mathbb{N}$. We denote the space of continuous and piecewise linear trajectories $h$ from $0,\infty)$ to $[0,1]$ such that $h(t)\in[0,1]$ for all $t\geq 0$ by $\mathcal{C}_{\rm PL}$. Each $h\in\mathcal{C}_{\rm PL}$ has at time $t$ a height $h(t)$ and a (right) slope
@@
Assume that for some $n\in\mathbb{N}_{0}\cup\{\infty\}$ we are given a configuration of pairs $(t_{i},a_{i})$ consisting of a birth time $t_{i}$ and an initial slope $a_{i}$
@@
[…]
Definition 2.1 (System of interacting trajectories, [16]).
[…]
@@
result from the following deterministic interactive dynamics on $(\mathcal{C}_{\rm PL})^{[n]\cup\{0\}}$, where we write $v_{i}:=v_{h_{i}}$ for $0\leq i<n+1$ in order to simplify notation:
  • One initial trajectory $h_{0}\in\mathcal{C}_{\rm PL}$ starts at height $h_{0}(0)=1$ with slope $v_{0}(0)=0$. All other trajectories $h_{i}$, $1\leq i<n+1$ start at height $h_{i}(0)=0$ with slope $v_{i}(0)=0$.
@@
  • For $i\geq 1$, at the birth time $t_{i}$ the slope $v_{i}$ of trajectory $h_{i}$ jumps from 0 to $a_{i}$.
@@
    $v^{\ast}:=\max\limits\{v_{i}(t-)\mid 0\leq i<n+1\mbox{ such that }h_{i}(t)=1\},$
    i.e. for all $0\leq i<n+1$ with $h_{i}(t)>0$
    $v_{i}(t):=v_{i}(t-)-v^{\ast}.$ (4)
  • Whenever a trajectory at some time $t$ reaches height 0 from above, its slope is instantly set to 0, and this trajectory then stays at height 0 forever.
    Let $e_{i}$ denote the time when this happens to the $i$-th trajectory during the sequential application of the heuristics (where we set $e_{i}=\infty$ if it never happens) where $0\leq i<n+1$. For $t\geq 0$ we say that the $i$-th trajectory is alive at time $t$ if $t_{i}\leq t<e_{i}$.
[…]
Definition 2.2 (Resident change times, resident type, and resident fitness, [16]).
Let $\mathbb{H}$ be as in (3), following the dynamics specified in Definition 2.1.
@@
    $\varrho(t)=\varrho^{(\beth)}(t):=\operatorname*{arg\,max}_{i}\{v_{i}(t-)\mid 0\leq i<n+1,h_{i}(t)=1\},$ (5)
    the resident type at time $t$, and we put $\varrho(0)=\varrho^{(\beth)}(0)=0$.
@@
The following remark is an adaptation of [16, Equations (3.5) and (3.6)] to our situation.
Remark 2.3 ([16]).
@@
    $f(t)=a_{\varrho(t)}+f(t_{\varrho(t)}),\quad t\geq 0$ (7)
    with $a_{0}:=0$.
@@
    (see [16, Remark 3.4, third arXiv version] for an explanation; this remark is not present in the current, fifth version of [16] on arXiv.).
Remark 2.4.
@@
At time 0, all $N$ individuals belong to the initial resident population and have the same fitness value 0. Beneficial mutations occur in a randomly chosen individual at time $t_{i}\log N$, with a positive fitness advantage (selective advantage) $a_{i}$, for all $1\leq i<n+1$. This selective advantage allows mutant subpopulations to reproduce faster; its strength does not depend on $N$ (this is why selection is called strong). Mutant subpopulations tend to either go extinct rapidly or grow approximately exponentially, hence their logarithmic population size divided by $\log N$ grows approximately linearly and takes $\Theta(\log N)$ time to reach the vicinity of 1.
@@
Speeding up time by a factor of $\log N$, we obtain approximately the system $\mathbb{H}(\beth)$; the corresponding convergence result is [16, Theorem 2.7]. Each mutation happens in an individual of the current resident population with high probability. Hence, the limiting trajectory $h_{i}$ starts with slope $a_{i}$ at time $t_{i}$ and keeps this slope until some subpopulation becomes resident (i.e. its trajectory reaches height 1). Upon this resident change, the slope of each subpopulation currently at a positive height drops by the increment of the resident fitness, i.e. the ultimate slope of the trajectory of the new resident before reaching height 1 (cf. (4)). Thus, slopes may become negative, corresponding to an approximately exponential decay of subpopulations that are already unfit compared to the current resident. The time when a trajectory hits zero marks the extinction time of the corresponding mutant subpopulation in the Moran model.
[…]
@@
In this case, we call $\mathbb{H}(\beth)$ the system of Poissonian interacting trajectories (or PIT for short) with parameters $(\lambda,\gamma)$, or briefly as the $\mathrm{PIT}(\lambda,\gamma)$. The corresponding trajectories and their (right) slopes will also be denoted by capital letters $H_{i}$ resp. $V_{i}$ (note that they are deterministic functions of the random input $((T_{j},A_{j}))_{j\in\mathbb{N}}$). The resident fitness $F$ that we mentioned in the introduction corresponds to $f$ from Definition 5 in the case of the PIT. The resident type at time $t\geq 0$ in this context will still be denoted by $\varrho(t)$. We will generally exclude conditions that are degenerate according to the following definition.
Definition 2.5.
[…]
@@
Claim 2.6.
Let $\beth=((T_{i},A_{i}))_{i\in\mathbb{N}}$ as above. Then we have that $\mathbb{P}(\beth\text{ is nondegenerate})=1.$
Remark 2.7.
@@
Remark 2.8 (The continued lines representation).
An alternative, equivalent representation of the PIT, called the continued lines representation, was provided in [16, first arXiv version, Section 5.3]. (The latter section is only included in the first and second version of [16] on arXiv.) This representation can be simulated without any calculations involving already decaying height functions, which turns out to be useful for algorithmic investigations. We describe this representation for a general system of interacting trajectories $\mathbb{H}(\beth)$ as follows.
[…]
Let us assume that $\mathbb{H}(\beth)$ is nondegenerate. Then, for $k\geq 0$, if $j\in\{0,1,\ldots,k\}$ satisfies $l^{r}(t_{k+1})=l_{j}(t_{k+1})$, we say that $j$ is the parent of $k+1$ (note that there is a unique such $j$ that is well-defined under the assumption of nondegeneracy).